OpenStudy (anonymous):

x divided by 5 equals 7 what does x equal to

OpenStudy (anonymous):

35

OpenStudy (one098):

x/5=7 Instead multiply 7 times 5 and that's your answer.

OpenStudy (anonymous):

x/5=7 solve this, you will get 5*7=35

OpenStudy (anonymous):

any problems, such as equations or inequalities, that have fractions and you do not want to mess with fractions...multiply everything by the common denominator and the fractions go away

OpenStudy (anonymous):

thanks @dtuniyants @ardentpikachu @One098 @kelliegirl33

OpenStudy (anonymous):

thats what we are here for :)

OpenStudy (one098):

^^ You're welcome!

OpenStudy (anonymous):

x+x+x=18 x=?

OpenStudy (anonymous):

@One098 @kelliegirl33

OpenStudy (anonymous):

6 6+6+6=18